# Using Large Multimodal Models to Extract Knowledge Components for Knowledge Tracing from Multimedia Question Information

### v1

@@ 3. Clustering Knowledge Components
Knowledge Component Quality Evaluation
The KCs used in this validation were evaluated by measuring the silhouette score of K-means clustering from 2 to 200 clusters. We automatically determined the optimal number of clusters by detecting the knee point of the silhouette score change (Rousseeuw 1987; Satopaa et al. 2011). To further verify the quality of KCs based on the level of clustering, we divided the entire range of 2 to 200 clusters into ten sections. For each section, we identified the point where the silhouette score was at its local maximum and observed the performance change of the AFM when KCs were generated with the corresponding number of clusters.
RMSE was measured not only for the entire dataset but also in environments where student ID information and item ID information were masked, respectively. This ablation experiment was done to identify which biases the model was utilized to demonstrate performance. For performance measurement, we utilized PyAFM, a Python implementation of the AFM (MacLellan, Liu, and Koedinger 2015).
To better understand the Knowledge Components (KCs) categorized through clustering, the most frequent name within each cluster was selected as the representative name for that cluster, alongside the descriptions included in that cluster. As illustrated in Box 1, these names typically consist of 2-3 words, such as Trigonometric Relationships or Moment Calculation. In most clusters, approximately one-third of the items shared the same name, which was then chosen as the mode and designated as the cluster’s representative name.
Knowledge Tracing Baselines Using KCs
With the extracted knowledge components, we utilized the code from prior works (Gervet et al. 2020b) to perform KT using IRT, PFA, DAS3H, SAKT, and DKT (Gervet et al. 2020a). In the experiments of the previous section, the performance of the OpenAI embedding model was superior to that of the T5-xxl model, so we evaluated the KCs generated using the OpenAI text-embedding3-large model. Among these, IRT is an algorithm that does not use KCs, while PFA and DAS3H use a Q-matrix. SAKT and DKT operate in environments with disjoint KCs, so we used joint skill assignments. Joint skill considers the combination of KCs assigned to each question as a single KC (Xiong et al. 2016). For example, a problem with KC 1 and KC 2 is treated as having a distinct KC different from a problem with only KC 1.
Results
Knowledge Component | Model | French | Computing | Statics | Biology | Psychology |
None | IRT | 0.822 | 0.809 | 0.797 | 0.743 | 0.781 |
PFA | 0.619 | 0.604 | 0.600 | 0.595 | 0.590 | |
Random | DAS3H | 0.873 | 0.816 | 0.804 | 0.762 | 0.793 |
SAKT | 0.828 | 0.780 | 0.812 | 0.858 | 0.809 | |
DKT | 0.925 | 0.817 | 0.860 | 0.912 | 0.822 | |
PFA | 0.787 | 0.723 | 0.751 | 0.666 | 0.698 | |
LLM (Ours) | DAS3H | 0.881 | 0.800 | 0.836 | 0.772 | 0.802 |
SAKT | 0.869 | 0.802 | 0.854 | 0.869 | 0.815 | |
DKT | 0.918 | 0.835 | 0.883 | 0.915 | 0.828 | |
PFA | 0.752 | 0.699 | 0.693 | 0.671 | 0.698 | |
Human | DAS3H | 0.911 | 0.840 | 0.843 | 0.768 | 0.801 |
SAKT | 0.850 | 0.554 | 0.854 | 0.874 | 0.817 | |
DKT | 0.929 | 0.868 | 0.877 | 0.918 | 0.828 |
@@ Results
Quality of the Generated Knowledge Components
Box 1 provides an example of generated KCs descriptions classified into the same cluster, represented as Trigonometric Relationships. Figure 3 shows the performance change of AFM with varying cluster numbers. When item information is masked, performance relies solely on the general ability of each student and KCs information, making it highly dependent on KCs quality. While overall and student-masked performance improved with more clusters, item-masked performance deteriorated, indicating that more clusters do not necessarily mean better KCs and suggesting room for improvement in verifying KCs consistency.
Table 3 compares the AFM performance of KCs selected using OpenAI’s embedding API, sentence-t5-xxl embeddings, and human experts in the statics, french, and computing domains. The results using OpenAI’s embeddings consistently outperformed sentence-t5-xxl. Given that previous research has shown the superior performance of OpenAI’s embedding models (Cao 2024), it can be concluded that as embedding models improve, the performance of KCs is likely to improve as well. However, both methods still showed comparable or better overall performance than human-created KCs, while item-blocked cross-validation performance was worse.
@@
Effect on Knowledge Tracing Performance
Table 2 shows the results of the KT experiments. For PFA and DAS3H, which are logistic regression-based KT models that can utilize multiple KCs (Pavlik, Cen, and Koedinger 2009; Choffin et al. 2019), we find using the KCs generated by our algorithm improves performance compared to the Random baseline. Notably, in PFA performance, our KCs outperform those of human experts across three domains. We believe that these advantages stem from the rich information provided by the multiple KCs tags per item.
@@
Towards More Refined Domain Modeling
@@
As reviewed by Pelánek (2017), Modeling the teaching domain with KCs has the potential to achieve greater granularity and richer informational content by assigning multiple KCs to a single item, modeling hierarchical relationships among KCs, or capturing prerequisite structures. However, in practice, it has often been the case that KCs are directly derived from curricula classifications, leading to a one-to-one mapping between problems and KCs (Sun et al. 2019). Consequently, Transformer-based KT models, such as SAINT and SAKT, have been designed to utilize this straightforward mapping (Choi et al. 2020; Pandey and Karypis 2019). However, given that real-world problems typically require the interconnection of multiple KCs, this approach deviates significantly from the intended role of KCs as “units of cognitive function” within the original Knowledge-Learning-Instruction framework (Koedinger, Corbett, and Perfetti 2012).
@@
Moreover, recent studies have increasingly diverged from the original definition of KCs, treating them as mere metadata to be leveraged for enhancing KT performance. There have been instances where KCs have been used interchangeably with terms like knowledge concepts (Ai et al. 2019) or generalized into labels such as knowledge, tag, or skill (Sun et al. 2019). This shift risks undermining the connection between KCs and learning science fields, such as recommendation systems that rely on KCs to provide meaningful educational insights. Therefore, it is crucial to maintain the integrity of KCs, ensuring they are not reduced to simple metadata for performance enhancement purposes.
@@
As large language models continue to advance, enabling the successful execution of more complex cognitive tasks, there is a growing opportunity to construct more sophisticated KC frameworks. These frameworks could consider hierarchical and prerequisite relationships with significantly reduced overhead, offering more robust and nuanced models for educational contexts. Moving beyond the AFM, there is a pressing need for enhanced methodologies that can leverage highly detailed KCs and evaluate their quality. Revisiting discussions on Cognitive Diagnosis Models may also help us remain focused on the core issues at hand (De La Torre 2011).
Linking to Zero-Shot Knowledge Tracing
@@
A key direction for future research, as proposed by this study, is the development of Zero-Shot KT techniques mediated by KCs. Currently, KT models are predominantly optimized for interaction data, rendering assessment infeasible without prior records. This limitation is particularly problematic for ITS, where KT models a student’s knowledge state based on their problem-solving history to predict future performance (Abdelrahman, Wang, and Nunes 2023). Despite the advances brought about by transformer architectures, these models still rely heavily on the statistical properties derived from problem-solving records. This approach contrasts sharply with human educators, who can intuitively assess knowledge and identify deficiencies without the need for extensive data histories.
@@
Biases stemming from an over-reliance on statistical data, which can be influenced by the difficulty level of the content or peer interactions.
@@
In the case of Deep Neural Network models, operations that are theoretically expected are not always guaranteed. For example, even if a learner answers more questions correctly than before, the learner’s KT prediction value may be lower than it was before answering those questions (Kim et al. 2021).
@@
The most significant synergy between automatically generated KCs and Zero-Shot KT lies in the potential to enhance KC frameworks without the constant need for human expert evaluation. By leveraging performance comparisons across existing KT benchmarks, much like the automated KC evaluation platforms provided by Datashop based on the AFM (Koedinger et al. 2010), researchers can accelerate the research cycle, eliminating human bottlenecks and fostering rapid advancements in educational technology.
Limitation
@@
The limitations of our study primarily stem from significant data losses during preprocessing, including the extraction of images from outdated HTML files, potential errors in converting audio to text, and the exclusion of content with uncertain mappings, which reduced transaction data by more than half. Additionally, current KT methodologies like SAKT, DKT, and AFM, which employ joint skill modeling, are not advanced enough to fully evaluate the effectiveness of our approach that assigns multiple KCs to each problem. Future research will need to focus on refining KT methodologies to better assess these detailed KCs. Lastly, as in the case of Fig2, we had to use images due to the nature of content where problem-solving heavily relies on visuals, and thus we were unable to conduct experiments comparing the performance differences between the Large Language Model and Large Multimodal Model.
Conclusion
In this study, we have proposed a novel approach that leverages instruction-tuned large multimodal models (LMMs) for the extraction and utilization of KCs from educational multimedia content. Our findings demonstrate that our automatically generated KCs can significantly enhance the performance of KT methods, sometimes even surpassing those of human experts.
To encourage further research in this promising area, we have released refined KT benchmarks across five domains. These benchmarks aim to facilitate the development of more advanced KT methodologies that can fully exploit the potential of automatically generated KCs, ultimately contributing to more personalized and scalable educational systems.
Appendix A Appendix
Data preprocessing
This section explains how the OLI datasets from CMU Datashop 666https://pslcdatashop.web.cmu.edu/ were preprocessed.
All transaction data was extracted by selecting “all data” in the “Export” tab and choosing “By Transaction” in the detailed options. The content data was collectively downloaded by clicking the “Download Problem Content” button in the “Dataset Info/Problem List” tab. We used the swftools package 777http://www.swftools.org/ in the Windows OS environment to extract images from the SWF file, which is not supported anymore due to the deprecation of flash support.
all mp3 files are converted into text files, using whisper-large-v2 (Radford et al. 2022).
OLI Engineering Statics
For the Statics dataset, we used the Fall 2011 version888https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=507. While other versions included a wider variety of KC models, the number of users and transactions was smaller compared to other statics benchmarks. Therefore, we considered the Fall 2011 version to be the superset and used it. The KC model named F2011 was used as the human tag. The original 361,092 transactions were reduced by about half to 189,047.
OLI Principles of Computing
The content of the OLI computing dataset 999https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1806 was different from other subjects, as it contained content from multiple subjects under the root directory. To maintain consistency with the method used for processing other datasets, only the content from Principles_of_Computing_v_1_10, which includes the version in the subject name, was used. The KC models principles_of_computing_1_13 and principles_of_computing_1_10 were used to calculate Table 3. The transaction data was filtered to include only those containing this content. As a result, the original transaction dataset, which had 37,233 rows, was reduced to 16,951 rows.
OLI French
We used the ’French1 - Spring 2014’ dataset for the French dataset101010https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=918. Since it included various KC models, we used Bonnie’s Model, which had the best AFM performance. For Table 3, Level4 KC model, which also reports good AFM performance, is used. This dataset contained many items with voice mp3 files, so we converted these files to text using the whisper-large-v2 model and inserted the transcriptions into the question text. At the position in the HTML where the mp3 execution was embedded, we prefixed the converted text with ’[transcription of embedded mp3 file]:’. Out of the total 278,489 rows of transactions, the final remaining data consisted of 53,255 rows.
OLI Biology
We used the ’Oli_biology’ dataset from the ’Bridge to Success’ project for the Biology data111111https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1148. The KC model named intro_biology-1_0 was used as the human label. After processing, 3,285,685 rows remained out of the original 5,852,795 rows of transactions.
OLI Psychology
We used the ’Psychology MOOC GT - Spring 2013’ dataset for the Psychology data121212https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=863 and utilized the KC model named psychology-1-4. After processing, 1,935,496 rows remained out of the original 2,493,609 rows of transactions.
Qualitative aspects of generated KCs
Tables 4, 6, 8, 10, and 12 present randomly selected content from each dataset, showing the GPT-4o generated names and descriptions for the KCs tagged to those contents, and the KC names classified by clustering. For the Biology and Psychology datasets, the higher specificity of the topics often resulted in all KCs within a single problem being classified under the same tag, explaining the relatively low Multi_KC ratio in Table 1.
Tables 5, 7, 9, 11, and 13 display 10 randomly selected descriptions from the first and second clusters of each dataset. Each cluster_name is determined by the most frequent KC_name among the KCs within that cluster as generated by GPT-4o, which may sometimes fail to accurately describe the cluster.
Detailed experiment settings
GPT-4o prompt
@@ Appendix A Appendix
Clustering
Figure 5 shows the WCSS and Silhouette Score of the K-means clustering method as the number of clusters varies from 2 to 200. Due to the instability of the Silhouette Score when the number of clusters is very small, we analyzed cases where the number of clusters is greater than 10. Each point in Figure 3 represents the local maximum Silhouette Score within one of the 10 bins divided between the range of 10 to 200 clusters, and the AFM performance was measured using these cluster numbers.
Meanwhile, for zero-shot implementation, using more than 100 KCs, which maximizes the Silhouette Score, made it impossible to have at least one problem with each KC in both the train and test splits. Therefore, we used the number of clusters at the local maximum of the third bin mentioned earlier, where such splits were feasible for all datasets. The selected numbers of clusters were 52, 63, 52, 61, and 49 for computing, statics, French, psychology, and biology, respectively.
Zero-shot Knowledge Tracing
We implemented zero-shot knowledge tracing using the same codebase131313https://github.com/theophilegervet/learner-performance-prediction that was used for KT implementation(Gervet et al. 2020a). The logistic regression setup was configured to experiment with various combinations of features, specifically utilizing the s, sc, tc, tw, w, and a tags. These options allow us to use features that record which KCs are present, how many times each KC appears in the user’s history, the total number of problems the user has solved, the total number of problems the user has correctly solved, and how many attempts were made for each problem within a time window. When selecting these options, we understood the code and chose all relevant options, but did not compare with other options due to the absence of a validation set.
AFM analysis
For AFM inference, we used the Python implementation compatible with the CMU DataShop format141414https://github.com/cmaclell/pyAFM. The mapped KCs were post-processed in the same scheme as when transactions are exported with the ‘By Transaction’ option from CMU DataShop, and the inference was performed using this code. We were unable to conduct AFM analysis on the Psychology and Biology datasets because the memory usage of the code increases proportionally with the transaction length. The required memory exceeded 40GB, which could not be handled by the computational resources used in this study, specifically a MacBook M3 Pro with 36GB of shared memory.
Generated Name | Description | KC Name |
Item ID: 663 | ||
call instruction | Understanding how the call instruction works. | function call |
extended stack pointer | Understanding what the extended stack pointer (ESP) is. | ESP register function |
ESP modification | Knowing that the call instruction modifies the ESP. | ESP register function |
Item ID: 317 | ||
Understanding Indexing | Knowledge of how to access elements in a list using their index values. | Index-based access |
List of Lists Structure | Understanding that elements of the main list can themselves be lists and how indexing works with nested lists. | Index-based access |
Cluster: signed char promotion |
Understand how and why an unsigned char is promoted to an unsigned int in an arithmetic expression involving only unsigned char operands. |
Know how multiplication of signed char values affects the resulting type in C. |
Knowing the C standard rules for promoting char to int during arithmetic. |
Knowing that arithmetic operations on smaller types (like signed char) promote them to int. |
Understand that signed char will be promoted to signed int when used in an arithmetic operation. |
Understand how operations on small integer types like char promote to int. |
Know that arithmetic operations on ’signed char’ operands promote them to ’int’. |
Understand how arithmetic operations affect signed char types and their promotion. |
Understand how unsigned char types are converted to int during type promotion. |
Understand how arithmetic operations affect signed char types and their promotion. |
Cluster: list iteration |
Knowing how to iterate over a list using a for loop. |
Using loops to iterate through the elements and indices of lists. |
Understand how to use len() to get the length of a list. |
Understanding how to iterate over elements in a list. |
Understanding how to iterate through elements of a list using a for loop. |
Understanding how to iterate through list elements using indexing. |
Understand how to use nested loops to combine two lists by iterating over them simultaneously. |
Understanding how to iterate through a list using indexing. |
Understand how to iterate through a list in Python. |
Iterate through the values of the integer list. |
Generated Name | Description | KC Name |
Item ID: 1273 | ||
Reading comprehension | Understand the text of the question and the options provided. | Pythagorean theorem |
Multiple choice format | Recognize the structure of a multiple choice question and how to select an answer. | Selecting correct option |
Decision making | Decide between the given options based on the question’s requirements. | Pythagorean theorem |
Item ID: 1097 | ||
Summation of Forces | Understand that F y = 0 denotes the summation of all forces in the y-direction equaling zero. | multiple choice |
Force Equilibrium | Recognize that the condition F y = 0 implies a state of force equilibrium in the vertical direction. | multiple choice |
Multiple Choice Questions | Know how to interpret and answer multiple-choice questions. | equilibrium conditions |
Selecting Correct Answer | Identify the correct option based on given conditions and context. | Pythagorean theorem |
Cluster: understanding question context |
Interpreting the context or scenario presented in the question. |
Identify the context and requirements of the given problem. |
Using context or surrounding information to make decisions or draw conclusions. |
Comprehend the context and specifics of the question to make an informed choice. |
Interpreting the question stem in part B to comprehend what is being asked. |
Comprehend the context and specifics of the question being asked to ascertain the validity of provided answers. |
Understanding the specific context of the problem to apply the correct principles. |
Identifying when a response should be definite or conditional based on given information. |
Interpreting the question and the content between the delimiters. |
Interpret the context to determine an appropriate logical statement. |
Cluster: equilibrium condition |
Understanding that all loads can be modeled and analyzed within two-dimensional space. |
Assess the position and arrangement of logs in a visual setup. |
Utilize the geometry of points and lines to find directional cosines. |
Understanding what ’as assumed’ and ’opposite to the assumed’ mean. |
CM is likely an abbreviation or notation that needs to be understood in context. |
Know when to use cos or sin based on the angle’s relation to the coordinate axes. |
using geometric relationships for problem-solving. |
Determine whether specific logs are touching or interacting. |
Determine whether one log is in contact with another. |
Understanding what ’as assumed’ and ’opposite to the assumed’ mean. |
Generated Name | Description | KC Name |
Item ID: 4237 | ||
Understanding French Vocabulary | Ability to comprehend and translate French words to English. | Personal pronouns in French |
Definition Matching | Ability to select the correct definition of a given word from a set of options. | Contextual understanding |
Contextual Knowledge | Understanding the context of common places and their descriptions, such as different types of stores. | subject-verb agreement |
Item ID: 6153 | ||
French Vocabulary | Understanding the meaning of French phrases and expressions. | imperative mood |
Phrase Matching | Identifying the correct definition for a given phrase. | French Greetings |
Cluster: French verb conjugation |
Translating French options to English to choose the correct one. |
Selecting the correct translation from given options. |
Selecting the appropriate translation from given multiple-choice options. |
Approach to verify the translated sentence through comparison with given options. |
Choosing the correct option from a list based on translation. |
Selecting the correct translation from given options. |
Identify the most contextually appropriate translation among multiple choices. |
Selecting the accurate translation option from multiple choices. |
Select the correct English translation from a list of options. |
Selecting the most suitable translation from multiple given choices. |
Cluster: negation |
Knowing the meaning and usage of the French word ’main(s)’. |
Understanding the correct English translation for the French word ’souper’. |
Understanding the French term ’être au régime’. |
Understanding how to translate interrogative phrases like ’Et les tiens?’ to ’What about yours?’. |
Understand the meaning of the French word ’parfois’. |
Understanding the phrase ’Vous avez de la chance.’ in French. |
Understanding how to translate ’moins le quart’ from French to English. |
Understand the meaning of the French phrase ’être en forme’. |
Understand the meaning of the French phrase ’faire la bise’. |
Understand the French phrase ’être en forme’. |
Generated Name | Description | KC Name |
Item ID: 1594 | ||
define allopatric speciation | the formation of new species in populations that are geographically isolated | definition of solute |
identify geographic barriers | understand that physical barriers can isolate populations, leading to new species formation | definition of solute |
differentiate speciation types | recognize the differences between allopatric and sympatric speciation | definition of solute |
Item ID: 89 | ||
capillary action in plants | Understanding how water moves up trees through capillary action caused by polar properties. | Bleach chemical properties |
polarity of water | Knowledge that water molecules are polar and that this property is essential for their movement through plant xylem. | Bleach chemical properties |
role of gravity in water movement | Recognizing that gravity is not the primary force in the upward movement of water in trees; instead, capillary action and transpiration play major roles. | DNA sequence function |
interpretation of true statements | Ability to evaluate multiple statements and identify the correct one based on knowledge of plant physiology and water properties. | pattern recognition |
Cluster: Catabolic pathways |
NADH is produced during cellular metabolism, not necessarily through direct energy creation by cells. |
Understanding that an enzyme is a type of protein. |
Understand that each metabolic pathway has key enzymes that catalyze major reactions. |
Knowledge that anabolic processes involve the building of complex molecules from simpler ones. |
Recognize that pyruvate is the main end product of glycolysis. |
Understanding whether glycolysis is dependent or independent of the electron transport chain. |
Identifying that the primary source of ATP in aerobic organisms is through cellular respiration, not photosynthesis, catabolic reactions, or anabolic reactions. |
A process in humans and other organisms where cells convert glucose and oxygen into energy (ATP), carbon dioxide, and water. |
Recognition that catabolic processes involve the breakdown or degradation of molecules. |
Understand the function and nature of enzymes in biological systems. |
Cluster: DNA sequence function |
Recognize that increasing temperatures can lead to sea level rise. |
Understanding the scientific explanations for the causes of natural disasters. |
Knowledge of how nutrients support the growth of algae in water bodies. |
Recognizing measures that reduce or prevent greenhouse gas emissions. |
Recognize what strategies effectively address or mitigate climate change. |
Understand the different types of greenhouse gases emitted by human activities. |
Understand how changes in greenhouse gas levels contribute to global warming. |
Understanding the role of plants in releasing water vapor into the atmosphere. |
Human-induced changes to the climate, primarily through the release of greenhouse gases. |
Recognize evaporation and condensation as key physical processes in various cyclical systems. |
Generated Name | Description | KC Name |
Item ID: 2066 | ||
Human Eye Anatomy | Understanding the different parts of the human eye and their functions. | Neurotransmitter identification |
Iris Location | The iris is positioned between the cornea and lens, controlling the amount of light that enters the eye. | Neurotransmitter identification |
Item ID: 709 | ||
Reasonable Mind Concept | Understanding that the decision involves logical planning and time management. | Pattern of Knowledge |
Wise Mind Concept | Recognizing the balance between emotional and reasonable mindsets, though not applicable here. | Role of Amygdala |
Emotional Mind Concept | Understanding that decisions driven purely by emotions are not being considered in this scenario. | Role of Amygdala |
Time Management Skills | The ability to plan and allocate time effectively for various activities including work, study, and social events. | Pattern of Knowledge |
Cluster: Role of Amygdala |
Recognizing that repeated rejection leads to intensified hurt feelings and negative emotional responses in males. |
Understanding that the amygdala is involved in processing emotions. |
Recognizing Donald’s physiological response (heart beating quickly) triggered by an external stimulus. |
Comprehending emotional language and appreciating complicated relationships among emotions. |
Determining which branch of emotional intelligence is being assessed by evaluating the context provided. |
Recognizing decisions driven primarily by emotions, such as impulsive spending. |
Recognize how emotional responses manifest due to minor, unexpected events. |
Secondary emotions are classified based on their valence (pleasant or unpleasant) and intensity (mild or intense). |
Understanding that actions taken in an emotional mind can lead to negative outcomes, like damaging a piece of furniture. |
Recognizing that the term ‘affective forecasting’ specifically pertains to predicting future happiness or emotional states. |
Cluster: option selection |
Ability to determine when symptoms began to occur. |
Recognizing the symptoms and effects of elevated cortisol levels, such as fatigue and depression. |
Recognizing compulsions such as repetitive hand-washing linked to fear of contamination. |
Recognize that fighting in combat is typically a traumatic experience. |
Identify symptoms such as avoidance of high places and intense fear related to heights. |
Knowing the symptoms specific to panic disorders. |
Knowledge of associating the onset of symptoms with specific days. |
Understand that dementia is not an inevitable part of the normal aging process. |
Recognize that the primary issues dealt with are depression, anxiety, and family stress. |
Understanding that comorbidity refers to an individual having two diagnoses at the same time, such as anxiety and depression. |

### v2

@@ 4 Results
Figure 4 provides an example of generated KCs descriptions classified into the same cluster, represented as Trigonometric Relationships. Figure 6 shows the performance change of AFM with varying cluster numbers. When item information is masked, performance relies solely on the general ability of each student and KCs information, making it highly dependent on KCs quality. While overall and student-masked performance improved with more clusters, item-masked performance deteriorated, indicating that more clusters do not necessarily mean better KCs and suggesting room for improvement in verifying KCs consistency.
@@
Table 3 compares the AFM performance of KCs selected using OpenAI’s embedding API, Sentence-T5-XXL embeddings, and human experts in the statics, french, and computing domains. The results using OpenAI’s embeddings consistently outperformed Sentence-T5-XXL. Given that previous research has shown the superior performance of OpenAI’s embedding models [3], it can be concluded that as embedding models improve, the performance of KCs is likely to improve as well. However, both methods still showed comparable or better overall performance than human-created KCs, while item-blocked cross-validation performance was worse.
@@
As qualitative aspects of generated KCs, Table 1(b) presents randomly selected content from the oli_statics and oli_psych-ology datasets, showing the GPT-4o generated names and descriptions for the KCs tagged to those contents and the KC names classified by clustering. For the Biology and Psychology datasets, the higher specificity of the topics often resulted in all KCs within a single problem being classified under the same tag, explaining the relatively low Multi_KC ratio in Table 2. Figure 4 displays example knowledge components belonging to two clusters in oli_statics dataset.
4.1 Effect on Knowledge Tracing Performance
| Knowledge Component Source | KT Model | French | Computing | Statics | Biology | Psychology |
| None | IRT | 0.822 | 0.809 | 0.797 | 0.743 | 0.781 |
| PFA | 0.619 | 0.604 | 0.600 | 0.595 | 0.590 | |
| Random | DAS3H | 0.873 | 0.816 | 0.804 | 0.762 | 0.793 |
| SAKT | 0.828 | 0.780 | 0.812 | 0.858 | 0.809 | |
| DKT | 0.925 | 0.817 | 0.860 | 0.912 | 0.822 | |
| PFA | 0.787 | 0.723 | 0.751 | 0.666 | 0.698 | |
| LLM (Ours) | DAS3H | 0.881 | 0.800 | 0.836 | 0.772 | 0.802 |
| SAKT | 0.869 | 0.802 | 0.854 | 0.869 | 0.815 | |
| DKT | 0.918 | 0.835 | 0.883 | 0.915 | 0.828 | |
| PFA | 0.752 | 0.699 | 0.693 | 0.671 | 0.698 | |
| Human | DAS3H | 0.911 | 0.840 | 0.843 | 0.768 | 0.801 |
| SAKT | 0.850 | 0.554 | 0.854 | 0.874 | 0.817 | |
| DKT | 0.929 | 0.868 | 0.877 | 0.918 | 0.828 |
Table 4 shows the results of the KT experiments. For PFA and DAS3H, which are logistic regression-based KT models that can utilize multiple KCs [33, 4], we find using the KCs generated by our algorithm improves performance compared to the Random baseline. Notably, in PFA performance, our KCs outperform those of human experts across three domains. We believe that these advantages stem from the rich information provided by the multiple KCs tags per item.
@@ 5 Discussion & Limitation
5.1 Towards More Refined Domain Modeling
@@
As reviewed by previous work [35], Modeling the teaching domain with KCs has the potential to achieve greater granularity and richer informational content by assigning multiple KCs to a single item, modeling hierarchical relationships among KCs, or capturing prerequisite structures. However, in practice, it has often been the case that KCs are directly derived from curricula classifications, leading to a one-to-one mapping between problems and KCs [42]. Consequently, Transformer-based KT models, such as SAINT and SAKT, have been designed to utilize this straightforward mapping [5, 32]. However, given that real-world problems typically require the interconnection of multiple KCs, this approach deviates significantly from the intended role of KCs as “units of cognitive function" within the original Knowledge-Learning-Instruction framework [19].
@@
Moreover, recent studies have increasingly diverged from the original definition of KCs, treating them as mere metadata to be leveraged for enhancing KT performance. There have been instances where KCs have been used interchangeably with terms like knowledge concepts [2] or generalized into labels such as knowledge, tag, or skill [42]. This shift risks undermining the connection between KCs and learning science fields, such as recommendation systems that rely on KCs to provide meaningful educational insights. Therefore, it is crucial to maintain the integrity of KCs, ensuring they are not reduced to simple metadata for performance enhancement purposes. We note that in this work we use "Knowledge Components" as our primary term and subsume related notions (e.g., concept, skill, tag) under this unified definition.
@@
As large language models continue to advance, enabling the successful execution of more complex cognitive tasks, there is a growing opportunity to construct more sophisticated KC frameworks. These frameworks could consider hierarchical and prerequisite relationships with significantly reduced overhead, offering more robust and nuanced models for educational contexts. Moving beyond the AFM, there is a pressing need for enhanced methodologies that can leverage highly detailed KCs and evaluate their quality. Revisiting discussions on Cognitive Diagnosis Models may also help us remain focused on the core issues at hand [10].
5.2 Linking to Zero-Shot Knowledge Tracing
@@
A key direction for future research, as proposed by this study, is the development of Zero-Shot KT techniques mediated by KCs. Currently, KT models are predominantly optimized for interaction data, rendering assessment infeasible without prior records. This limitation is particularly problematic for ITS, where KT models a student’s knowledge state based on their problem-solving history to predict future performance [1]. Despite the advances brought about by transformer architectures, these models still rely heavily on the statistical properties derived from problem-solving records. This approach contrasts sharply with human educators, who can intuitively assess knowledge and identify deficiencies without the need for extensive data histories.
@@
Biases stem from an over-reliance on statistical data, which can be influenced by the difficulty level of the content or peer interactions.
@@
In the case of Deep Neural Network models, operations that are theoretically expected are not always guaranteed. For example, even if a learner answers more questions correctly than before, the learner’s KT prediction value may be lower than it was before answering those questions [17].
@@
The most significant synergy between automatically generated KCs and Zero-Shot KT lies in the potential to enhance KC frameworks without the constant need for human expert evaluation. By leveraging performance comparisons across existing KT benchmarks, much like the automated KC evaluation platforms provided by Datashop based on the AFM [18], researchers can accelerate the research cycle, eliminating human bottlenecks and fostering rapid advancements in educational technology.
5.3 Limitations
@@
Our study has several limitations. First, our data preprocessing led to significant losses—extraction from outdated HTML (including deprecated Flash content), potential audio-to-text conversion errors, and the exclusion of uncertain mappings reduced transaction data by more than half. Second, current KT models (SAKT and DKT) that rely on joint skill modeling are not sufficiently advanced to fully assess our detailed multi-KC approach. Finally, due to the visual nature of some problems, we relied on images rather than text, which prevented direct comparisons between Large Language Models and Large Multimodal Models.
5.3.1 Losses in Data Preprocessing
Refining the content data proved challenging, resulting in substantial losses. Extracting images from HTML files including deprecated Flash elements and converting audio to text introduced errors. The most substantial data loss occurred during the mapping of content data to transaction data. Due to incomplete metadata and the absence of content for some problems, we excluded any content with uncertain mapping to ensure data accuracy. This process reduced the original transaction data by over half.
5.3.2 KT Model Limitations
SAKT and DKT rely on joint skill modeling, which does not fully capture the nuances of our multi-KC approach. Additionally, AFM—being an older model—utilizes relatively simple feature engineering to process time-series data, limiting its ability to evaluate advanced KC details. Our primary focus was on developing detailed KCs, leaving further refinement of KT methodologies for future work. Moreover, it would be desirable for future work to support additional KT frameworks—such as the multi-method pyKT library [24] and the LKT framework [34]—so that our extracted KCs can be evaluated across a broader and more diverse set of models.
5.3.3 Access to Raw Content Data
While we have made all generated tags and refined data publicly available, the parsed raw data remains accessible only through CMU DataShop due to their policy. Because these contents are actually used in educational settings, they cannot be made publicly available without restrictions. However, obtaining access is straightforward and promptly processed, and we have provided reproducible preprocessing code to facilitate this. We are committed to ensuring the usability of this benchmark by providing all necessary support for access.
5.3.4 Upper Bound on the Number of Clusters
As shown in Table 6, AFM’s overall RMSE continues to decrease even up to 200 clusters. In theory, a reversal would be expected — an increase in RMSE due to overfitting — when the cluster count becomes sufficiently large, but our experiments were restricted to at most 200 clusters. Moreover, we evaluated only three relatively small datasets (Statics, Computing, and French), which further limited our ability to explore higher cluster counts. These restrictions stem from the computational burden that grows with both the size of the dataset and the granularity of the cluster under finite resources. We anticipate that the development of more efficient algorithms for selecting the optimal number of KCs would enable exploration beyond this bound and could yield additional performance gains.
6 Conclusion
We have presented a novel, zero-shot approach that leverages instruction-tuned large multimodal models (LMMs) to automatically extract knowledge components (KCs) from educational multimedia content. Unlike traditional methods—which rely on human-generated labels or purely statistical techniques—our approach directly parses text, images, and audio to generate detailed KCs. Experimental evaluations across five domains and multiple knowledge tracing (KT) models (including IRT, PFA, DAS3H, SAKT, and DKT) demonstrate that the LMM-generated KCs not only match but often exceed the performance of human-defined KCs, thereby improving the accuracy and interpretability of student performance predictions.
In addition, by releasing refined KT benchmarks enriched with these automatically generated KCs, we provide a valuable resource for the community to further develop advanced KT methodologies. While our findings highlight the promise of automated KC extraction in enhancing intelligent tutoring systems, they also reveal key limitations—such as significant data losses during preprocessing and the constraints of existing KT models in fully capturing the nuances of multi-KC assignments—that must be addressed in future research. Moving forward, integrating more sophisticated domain modeling techniques and exploring zero-shot KT strategies will be crucial for developing more personalized and scalable educational systems.